Please help
1. A right angles triangle has sides in the ratio 3:5:7. If the hypotenuse is 84mm, find the length of the other two sides.
2. A mass is composed of 3 parts of copper and 4 parts of Zinc. Find the mass of copper and Zinc in a casting with a mass of 98kg.

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

duplicate.
-----
That's not a right triangle.

Answer by stanbon(75887)

1. A right triangle has sides in the ratio 3:5:7. If the hypotenuse is 84mm, find the length of the other two sides.
------------------
7x = 84
x = 12
-----
Sides::
3x = 36
5x = 60
7x = 84
--------------------------
2. A mass is composed of 3 parts of copper and 4 parts of Zinc. Find the mass of copper and Zinc in a casting with a mass of 98kg.
---------------------------
copper = 3x
zinc = 4x
----------------------
3x + 4x = 98 kg
7x = 98
x = 14
------
copper = 3x = 3*14 = 42 kg
zinc = 4x = 4*14 = 56 kg
------------
Cheers,
Stan H.
--------------